Status Report

NASA Space Station On-Orbit Status 17 Feb 2004

By SpaceRef Editor
February 17, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 17 Feb 2004

All ISS systems continue to function nominally, except as noted previously or below.

Onboard sleep cycle continues with a 9:00pm wakeup and 12:30pm bedtime.  Foale and Kaleri reported that they have well adapted to the 4-hr. shift and are quite happy with it.

The crew spent several hours in the DC-1 docking compartment on preparations of Orlan spacesuit equipment for the Soyuz ingress demonstration scheduled for 2/19 (Thursday).  Main focus was on EVA support systems configuration and checkout, communications setup, leak checks and valve functionality tests.  All tasks were satisfactorily completed. [For the duration of these activities, MCC-H reconfigured Space-to-Ground/audio capability to allow the crew to talk with the ground from inside the DC-1.  After Kaleri had checked out the EVA support panel (POV) in the DC-1 airlock, the crew activated and visually inspected the two Orlan-M suits #14 (blue) for Foale and #23 (red) for Kaleri, along with their BRTA radio telemetry units.  Wearing their medical instrumentation belts, they tested comm and biomedical telemetry via the BSS interface system for vital signs and equipment monitoring, with alarm signals from the BSS rerouted to the POV panels in the DC-1 and Service Module (SM).  Suit consumables (such as filters, LiOH cans and batteries) were installed and auxiliary equipment prepared and installed, such as the BK-3 oxygen (O2) tanks.  The suits were preliminarily fit-checked in size (height), the latter to be repeated under reduced airlock pressure.  The Orlans and their BSS interface control units were checked out for proper air/water separation of their cooling system and tested for leaks of their reserve bladders.  The activities were supported by tag-ups with ground specialists.]

The ground has successfully activated the Elektron oxygen generator, and it is now running nominally on 32 amps. [Background & ppO2 management: After all attempts at restoration of nominal operation failed in the last three days, Russian specialists changed the cycling of the buffer tank that adjusts the amount of condensate sent through the unit.  This in effect changed the hydraulic resistance of the water supply line, raising water pressure and thereby mitigating the machine’s sensitivity to the small pressure variations that caused the pump shutdowns.  Running now smoothly on the primary pump, the Elektron, according to TsUP/Moscow, appears to have been “cured”, at least for the moment.  In the 32A mode, it can raise cabin air ppO2 (oxygen partial pressure) by 1 mmHg per day.  To correct the currently reduced ppO2 of 155 mmHg, as measured by the MCA (Major Constituents Analyzer, today activated briefly for this purpose) along with total pressure, an O2 repress from Progress 13P of 5-8 mmHg is planned.  Should the Elektron fail again before the EVA-9, two represses, totaling 15-16 mmHg, would be required.]

The crew continued cargo transfers from the Progress, assisted by a detailed Russian & US transfer/unpack list and the IMS (inventory management system).

Mike Foale conducted the daily routine maintenance on the SOZh life support system, comprising the water supply equipment, food supply subsystem (SOP), and sanitary hygiene equipment (SGO). [Maintenance generally consists of inspection of fluid lines/connections and filter replacement.]

Alex Kaleri broke out the “Urolux” equipment and set it up for his and Mike’s next session with the Russian biochemical urinalysis test (PZE MO-9), on tomorrow’s schedule for both crewmembers.  Also scheduled for tomorrow for them is their second session with the MO-5 MedOps protocol of cardiovascular evaluation during graded exercises on the VELO cycle ergometer.

At 4:55am EST, Foale and Kaleri downlinked a live non-interactive TV message to the annual Houston Livestock Show and Rodeo, via Ku- and S-band.

The crew worked out according to their regular daily physical exercise program on TVIS treadmill and CEVIS cycle (aerobic), VELO ergometer with force loader and RED exerciser (anaerobic).

The 2.5-in. long FOD (foreign object debris) observed floating away from the ISS on 2/15 has been identified by Moscow, thanks to the ID number visible on it, as an eyebolt from the Progress 13P starboard solar array’s holddown mechanism.  There are four of these bolts on the Progress, and also on Soyuz, and it is not clear at this time how it came lose and if other pieces were also separated. [The eyebolts engage in hooks to keep the folded arrays fastened down during launch and ascent.  They are no longer needed after the arrays are released by pyrotechnic charge and deployed by spring force, but they should remain attached to the vehicle structure.]

To support the debris analysis, Sasha powered up the SSRMS (Space Station Remote Manipulator System) after wakeup last night at 9:00pm, for the ground to remotely control it overnight through a survey of the Progress and Soyuz solar array wing tips with the tip elbow (TE) and Base Elbow (BE) camera, as far as their field of view allowed. [SSRMS can view Soyuz’ port and starboard arrays, but only the port array on Progress, whose out-of-sight starboard array happens to be the FOD’s origin.]
The Russian SKV-2 air conditioner, with its new coolant line connected by the crew on 2/13, has not yet been tested. [After its installation, the new line, with its larger bend, was found to be in direct contact with station structure, which in the longer term could lead to its damage due to vibration.  Any further testing or other SKV-2 activity has been deferred until after the EVA-9.] 

SKV-1, on thermal loop #1 (KOV-1) continues to operate, but at an efficiency level which has increased the load on the U.S. CCAA dehumidifier (common cabin air assembly).  As a result, the Lab collection tank has currently accumulated about 72 lbs of condensate, approaching the tank limit of 100 lbs  which will require offloading of its contents to a CWC (collapsible water container).    [Offloading to CWC takes up about two hours crewtime (one person), on an already crammed timeline.  Current SKV-1/KOV-1 operation results in ~2 lbs of U.S. condensate per 24-hr. day, which rises to 6 lbs/day if SKV-1 is off.  Moscow was asked to run their SKV-1/KOV-1 system in maximum collection mode to help avoiding this crewtime “hit” prior to the EVA.]

On 2/19, as preparation for the EVA-9, MCC-H will conduct a test of the ability to send commands to the USOS (U.S. segment) through RS (Russian segment) assets.  Purpose of testing this contingency configuration is (a) to provide proficiency to ground controllers in the event it is required during the spacewalk, and (b) to acquire data on command-link “latency”, i.e., the time lag between command uplink and command execution, given that in a worst-case situation the command has to “queue up” on the RS server, loaded with interfering Russian signals.

Moscow reports that yesterday’s MCS (motion control system) test on Soyuz TMA-3/7S was successful.  It also certified the berthing thrusters for a redocking. [The test, on manifold #2 of the KDU integrated propulsion system, was also intended for checking the line with the helium leak between the He pressurization tanks and the prop tanks of the KDU, which had been observed during the flight of 7S on 10/18-20/2003.  The line, which had leaked down to zero since then, was successfully purged with He yesterday but now is again exhibiting its gradual pressure drop.  The test did not yield further insight into the root cause of the leak, which still allows functionality of the affected string 2, although manifold #1 will be used as primary string.  The latter remains ready for undocking & deorbit, and there are also sufficient He and prop for undocking & deorbit in the closed-off manifold #2.]

Last night’s repeat of the checkout of the BCC (backup control center), rescheduled from 2/12, was successfully completed.  Purpose of the periodic dry-run is to demonstrate BCC functionality and provide proficiency training for HSG (Houston Support Group) personnel at the TsUP/Moscow HSR (Houston Support Room). [The 2/12 BCC C/O was only partially successful and was not completed.  BCC utilizes the HSG command server and telemetry server.  In last night’s test mode, HSG (as BCC) sent a single not-for-uplink PPCP (preplanned command packet) to MCC-M/TsUP, then canceled it after successful transmission.  MCC-M played back ACT (American contingency telemetry) recorded earlier (11/19/03), with HSG confirming successful receipt.  TsUP commanding to the Russian segment (RS) via U.S. S-band as well as MCC-H-to-TsUP flow of status telemetry were also not affected, but MCC-H has no real-time telemetry and command capability via Russian ground stations (RGS) during these dry-runs.]

The crew’s sleep time began at 12:30pm EST, to extend until tonight, 9:00pm.

Today’s CEO targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, except for the shutter closure and condensation-prevention plan (limited to 90 min. in 24 hours), were Beijing, China (looking just left of track for the vast urban region of the North China plain–from the foothills of the plateau all the way to the sea.  This pass included Beijing and its major city of Tianjin), Internal waves, Philippines (the sunglint point passed over an area of interest between Mindanao, the southernmost big island of the Philippines, and Borneo.  Looking right between one and four degrees off track), Dhaka, Bangladesh (with a population close to 10 million, this is a mega city whose growth, much faster than the high national rate, is being watched.  Forty-four percent of Bangladesh’s urban population lives in Dhaka.  Two 180-mm frames may be sufficient to capture the urban region), Rangoon, Myanmar (nadir pass over this capital city of 4 million.  Due to political problems, this is one major Asian city that has lost population in the last ten years), Plankton bloom, Pakistan (Dynamic event.  The strongest blooms lie mainly along the coastlines of Pakistan and India [nadir and left]), Karachi, Pakistan (nadir pass over Pakistan’s largest city), Bombay, India (looking left of track on the major bay in the coastline), and Sea ice, James Bay (Dynamic event.  ISS passed across southernmost Hudson Bay.  The extent and distribution of ice, and its evolution through time, are highly dynamic).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of 3:00am EST, 2/13).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On (32A).  Vozdukh CO2 scrubber is On (in Automatic Mode).  U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode).  TCCS (trace contaminant control subsystem) is operating.  SM Gas Analyzer has been calibrated and is used for ppO2 and ppCO2 monitoring.  MCA (major constituents analyzer) is on to verify CSA-CP readings and support O2 repress (was in Life Extending Mode).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On, SKV-2 is Off (repair now completed; to be tested ASAP).
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 25.7; ppO2 (mmHg) — 137.6; ppCO2 (mmHg) — 2.5;
  • SM Transfer Compartment:  Pressure (mmHg) — 756; temperature (deg C) — 20.0.
  • FGB Cabin:  Pressure (mmHg) — 744; temperature (deg C) — 22.3.
  • Node:  Pressure (mmHg) — 735.62; temperature (deg C) — 23.2 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 737.95; temperature (deg C) — 24.8; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 738.15; temperature (deg C) — 24.0; shell heater temp (deg C) — 24.2, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1:  Shell heater temp (deg C) — 22.2
  • PMA-2:  Shell heater temp (deg C) — 10.5

(n/a = data not available)

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in Directed Position (2B: 235 deg; 4B: 125 deg); non-suntracking, “night glider”/”sun slicer” drag reduction mode.
  • SM batteries:  Battery #8, formerly known as #7, is failed (to be replaced); all other batteries (7) are in “Partial Charge” mode.
  • FGB batteries:  Battery #3 is off (capacity restoration mode, ROM); battery #2 is in “Cycle” mode; all other batteries (4) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22).

Propulsion System:

  • Total propellant load available: 4070 kg (8972 lb) as of 2/12/04  [SM(755) + FGB(2656) + Progress M(0) + Progress M-1(659)].  (Capability: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH YVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, y-axis in velocity vector [yaw: -90 deg, pitch: -9 deg, roll: 1.7 deg]), with CMG/TA (thruster assist) Momentum Management.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • *S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 7:26am EST [= epoch]):

  • Mean altitude — 366.8 km
  • Apogee — 371.4km
  • Perigee — 362.2 km
  • Period — 91.88 min.
  • Inclination (to Equator) —  51.628 deg
  • Eccentricity — 0.000682
  • Orbits per 24-hr. day — 15.67
  • Mean altitude loss last 24 hours — 100 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 29938

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see

SpaceRef staff editor.